OpenStudy (anonymous):

select the expression that represents a rational function: 7=6/3x-2 f(x)=6/3x+2 x/6=6/3x+2 y=3x+2 plz help

OpenStudy (mathmate):

@sam_avd brackets are critical for this question. I think you mean 7=6/(3x-2) f(x)=6/(3x+2) x/6=6/(3x+2) y=3x+2 Hint: Two of the choices are equations, not functions. Of the remaining two, we can distinguish a rational function by the definition that a rational function is a function of the form \(\large \frac{p(x)}{q(x)}\) where p(x) and q(x) are polynomials. see also http://www.math.utah.edu/~wortman/1050-text-rf.pdf

OpenStudy (anonymous):

thank you so much!;)